OpenStudy (anonymous):

Let cos Y =.98 what is CSC Y?

OpenStudy (raden):



OpenStudy (raden):

to get cscY, you have determine the length of x then use cscY = 100/x = 100/... = ?

OpenStudy (raden):

* have to

OpenStudy (raden):

use the pythagorean theorem : x^2 = 100^2 - 98^2 = (100+98)(100-98) = 198 * 2 = 396 x = sqrt(396) = sqrt(36 * 11) = sqrt(36) * sqrt(11) = 6sqrt(11) so, cscY = 100/x = 100/6sqrt(11)

OpenStudy (anonymous):

I got 5.05 not sure I m right

OpenStudy (raden):

if you want rationalize it, just multiply by sqrt(11)/sqrt(11) cscY = 100/6sqrt(11) * sqrt(11)/sqrt(11) = 100sqrt(11)/(6*11) = 100sqrt(11)/66